# Prospects for Cosmological Research with the FAST Array: 21-cm Intensity Mapping Survey Observation Strategies

### v1

@@ IV Conclusion
Secondly, we discuss the combination of the 40 m dishes and FAST. Incorporating 40 m dishes with FAST significantly enhances survey efficiency, thereby reducing BAO parameter observational errors to typically below 2%. For the combination with FAST, 6$\times$40 m dishes are a more balanced configuration compared to 20$\times$40 m dishes for 21 cm IM.
@@
Acknowledgements.
We thank Guang-Chen Sun, Ji-Guo Zhang, Zi-Qiang Zhao, and Lu Feng for helpful discussions and suggestions. We are grateful for the support from the National SKA Program of China (grant Nos. 2022SKA0110200 and 2022SKA0110203), the National Natural Science Foundation of China (grant Nos. 11975072, 11875102, and 11835009), and the National 111 Project (Grant No. B16009).

### v2

@@ 4 Conclusion
Secondly, we discuss the combination of the 40-m antennas and FAST. Incorporating 40-m antennas with FAST significantly enhances survey efficiency, thereby reducing BAO parameter errors to typically below 2%. For the combination with FAST, 6$\times$40-m antennas are a more balanced configuration compared to 24$\times$40-m antennas for 21 cm IM.
@@
Acknowledgments
We thank Guang-Chen Sun, Ji-Guo Zhang, Zi-Qiang Zhao, and Lu Feng for helpful discussions and suggestions. We are grateful for the support from the National SKA Program of China (Grants Nos. 2022SKA0110200 and 2022SKA0110203), the National Natural Science Foundation of China (Grants Nos. 12473001, 11975072, 11875102, 11835009, and 12305069), the National 111 Project (Grant No. B16009), and the Program of the Education Department of Liaoning Province (Grant No. JYTMS20231695).